• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Reitz

    Log entry time 07:52:46 on November18,2002

    Entry number 90235

    keyword=Owl Shift Summary


    shortly after 0.00 5uA beam arrived, did Harp measurement

    MCC called, they want to take beam away for 20 min, to work
    on the frequent beam trips

    shortly after 1.00 am 5uA beam
    1) chamber test, chamber positioned next to big bite

    1:20 we asked for controlled access to move chamber to second position

    2:40 we are done, MCC has problems with Moeller Target

    3:20 Moeller Target has still problems, we asked for controlled access
    to move wire chamber to most interesting position behind BigBite,
    since we are already behind schedule

    around 5:00 MCC finally fixed Moeller problem (after consulting with Eugene
    and entering the Hall they found a loose wire on an end switch)

    5:10 MCC called: vacuum faults in A line, several valves closed
    Target vacuum didnt show anything suspicious in that time (6*10**-6 Torr)

    8:00 wire chamber tests ended, Mike Seely takes over for target work